Circular No. 8690

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET C/2006 F1 (KOWALSKI) R. A. Kowalski reports the discovery of a comet on CCD images taken

in poor seeing (4′′ FWHM) with the 1.5-m reflector in the course of the Mt. Lemmon Survey (discovery observation tabulated below); the object appeared slightly condensed with a diameter of ≈ 8′′, elongated in p.a. ≈ p.a. 265^{o} (though no obvious tail). On images taken with the same telescope by E. J. Christensen on Mar. 23.5 UT, the object displays a condensed 8′′ coma and a 10′′ tail in p.a. ∼ 245^{o}. CCD images taken by P. Birtwhistle with a 0.40-m f/6 Schmidt-Cassegrain reflector at Great Shefford, U.K., on Mar. 23.1 show the object with a 8′′ coma and a tail 15′′ long in p.a. 245^{o}.

2006 UT  α_{2000}  δ_{2000}  Mag.
Mar. 21.49155  16^{h}21^{m}29.^{s}96  +2^{o}00′22′′.3  18.8

The available astrometry, the following parabolic orbital elements, and an ephemeris appear on MPEC 2006-F20.

T = 2007 May 19.709 TT  ω = 208.^{o}890
Ω = 124.985 }2000.0
q = 1.93896 AU  i = 21.585

SUPERNOVAE 2006at, 2006av, 2006aw Further to IAUC 8666, L. A. G. Monard reports his discovery of two

apparent supernovae on unfiltered CCD images:

SN  2006 UT  α_{2000}  δ_{2000}  Mag.  Offset
2006av Feb. 19.776 2^{h}57^{m}46.^{s}48 −36^{o}42′26′′.2 18.3 30′′ W, 38′′ N
2006aw Mar. 15.747 5 04 18.18 −63 34 57.3 14.9 7′′ W, 2′′ S

Additional approximate magnitudes from Monard for 2006av in ESO 356-22: Jan. 31.757 UT, [18.5; Feb. 28.80, 18.7 ± 0.4 (mediocre conditions; stacked images); Mar. 5.735, 19.0 ± 0.2; 20.774, 18.3 ± 0.3. Nothing is visible at the location of 2006av on the Digitized Sky Survey in all bands (including limiting red mag 20.5). Additional approximate magnitudes for 2006aw in ESO 85-G38: Feb. 16.783, [18.5; Mar. 15.858, 14.9 ± 0.2; 19.751, 15.2. Nothing is present at the location of 2006aw on a red Digitized Sky Survey image (limiting mag 20.5).

SN 2006at (cf. IAUC 8687) has been confirmed to be a type-II supernova by S. Blondin et al. (details on CBET 441).

2006 March 24  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2006 CBAT  Daniel W. E. Green